Olaf is standing on a sheet of ice that covers the football stadium parking lot in Buffalo, New York; there is negligible friction between his feet and the ice. A friend throws Olaf a ball of mass 0.400 kg that is traveling horizontally at 10.2 m/s . Olaf's mass is 73.5 kg . Olaf is standing on a sheet of ice that covers the football stadium parking lot in Buffalo, New York; there is negligible friction between his feet and the ice. A friend throws Olaf a ball of mass 0.400 kg that is traveling horizontally at 10.2 m/s . Olaf's mass is 73.5 kg

A) If Olaf catches the ball, with what speed

vf do Olaf and the ball move afterward?

B) If the ball hits Olaf and bounces off his chest horizontally at 7.90

m/s in the opposite direction, what is his speed vf after the collision?
Answer #1

GIVEN :ball of mass= 0.400 kg,speed=10.2 m/s, Olaf's mass is 73.5 kg

a)) If Olaf catches the ball, with what speed

vf do Olaf and the ball move afterward?

initial momentum = final momentum
(0.400 kg)(10.2 m/s) = (73.5 + 0.400 kg)(v_f)
v_f = +0.0552 m/s

B) If the ball hits Olaf and bounces off his chest horizontally at 7.90

m/s in the opposite direction, what is his speed vf after the collision?

(0.400 kg)(10.2 m/s) = (73.5 kg)(v_f) + (0.400 kg)(-7.90 m/s)
v_f = +0.0985m/s